# HTMformer: Hybrid Time and Multivariate Transformer for Time Series Forecasting

### v2

@@ Abstract
Transformer-based methods have achieved impressive results in time series forecasting. However, existing Transformers still exhibit limitations in sequence modeling as they tend to overemphasize temporal dependencies. This incurs additional computational overhead without yielding corresponding performance gains. We find that the performance of Transformers is highly dependent on the embedding method used to learn effective representations. To address this issue, we extract multivariate features to augment the effective information captured in the embedding layer, yielding multidimensional embeddings that convey richer and more meaningful sequence representations. These representations enable Transformer-based forecasters to better understand the series. Specifically, we introduce Hybrid Temporal and Multivariate Embeddings (HTME). The HTME extractor integrates a lightweight temporal feature extraction module with a carefully designed multivariate feature extraction module to provide complementary features, thereby achieving a balance between model complexity and performance. By combining HTME with the Transformer architecture, we present HTMformer, leveraging the enhanced feature extraction capability of the HTME extractor to build a lightweight forecaster. Experiments conducted on eight real-world datasets demonstrate that our approach outperforms existing baselines in both accuracy and efficiency.
@@ 1 Introduction
Long-term time series forecasting holds significant importance in various fields such as finance and economics (Andersen et al., 2006), climate science (Mudelsee, 2019), healthcare (Zeger et al., 2006), geophysics (Gubbins, 2004), industrial monitoring (Truong et al., 2022), etc. Recently, Transformer (Vaswani et al., 2017) and its variants have achieved tremendous success in time series forecasting, showing great promise. Current Transformer-based forecasters often overemphasize modeling temporal dependencies. However, the effective information within the temporal dimension is limited. Consequently, existing models incur increasing computational overhead without significant accuracy improvements.
Unlike natural language sequences, time series data, particularly in domains such as transportation, meteorology, and electricity systems, are influenced by specific physics-based systems, e.g., traffic flow is largely governed by the real-world road network. Therefore, multivariate correlations are essential for time series forecasting tasks (Zheng & Sun, 2024). Two key points are worth noting: 1) the vast majority of features in time series are contained within the temporal dimension, and the depth of exploration in this dimension significantly influences prediction accuracy (Hyndman & Athanasopoulos, 2021); 2) the multivariate dimension also contains a large number of features (Pravilovic et al., 2013), and fully exploiting multivariate features can further boost performance.
The performance of Transformer-based forecasters is highly dependent on the design of the embedding layer. Before applying multi-head self-attention to the embeddings, effective feature extraction is critical to enhancing the model’s sequence modeling capacity (Li et al., 2023a). We compare the ground-truth values and the predictions obtained under different embedding strategies on four datasets (see Appendix A). Specifically, models such as iTransformer (Liu et al., 2024b) and PatchTST (Nie et al., 2023), which employ an inverted embedding strategy or patch-based embedding strategy for deeper time-series representation learning, achieve significantly better performance than the vanilla Transformer, which only uses simple convolutional embeddings. Meanwhile, we also notice that as the input length increases, the prediction loss fails to markedly decrease and even tends to rise. These suggest that the modeling capacity of Transformer-based forecasters is inherently constrained by the quality of embeddings, and this limitation cannot be eliminated by increasing the lookback length.
@@
To mitigate this issue, this paper proposes to jointly capture temporal and multivariate features in the embedding layer, thereby yielding semantically rich embedding representations to overcome the inherent limitations of the conventional embedding strategy (see Appendix B). Utilizing proposed strategy, we constructed a Transformer-based forecaster and achieved state-of-the-art empirical performance on benchmarks, as shown in Figure 1. The contributions of this work are as follows:
  • We propose the Hybrid Temporal and Multivariate Embedding (HTME) extractor, which comprises a lightweight temporal feature extractor to capture informative temporal dynamics and a multivariate feature extractor to model the multivariate dependencies. The HTME is broadly applicable across Transformer-based forecasting architectures and effectively enhances the capacity of transformer-based models to model complex time series.
  • We introduce a novel forecasting model, HTMformer, to further validate the effectiveness of the HTME strategy. In this framework, the HTME serves as input to a vanilla encoder-only Transformer framework. Extensive empirical evaluations are conducted on eight real-world benchmarks spanning application domains. Notably, HTMformer consistently achieves state-of-the-art performance across various benchmarks.
@@
Transformer-Based Time Series Forecasting.
Well-designed Transformer-based forecasters have achieved substantial success in time series forecasting tasks. They are not limited by the inability of CNN-based methods (Liu et al., 2022) to capture long-term temporal dependencies (Sutskever et al., 2014; Karim et al., 2017), nor do they encounter the vanishing and exploding gradient problems (Bengio et al., 1994; Hochreiter & Schmidhuber, 2019) commonly observed in RNN-based methods (Cho et al., 2014; Lai et al., 2018; Salinas et al., 2020). In recent years, advanced models such as Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022), and Crossformer (Zhang & Yan, 2023) have primarily introduced novel attention mechanisms and architectural modifications for explicitly modeling long-term dependencies and temporal interactions in time series data. Nonetheless, as discussed in the Introduction, Transformer-based forecasters suffer from inherent limitations in effectively modeling time series. Refining the attention mechanism and architecture alone cannot address this limitation. Unlike prior works, we propose a novel embedding strategy for comprehensive feature extraction on the raw input before it is fed into the attention module, thereby substantially improving the modeling capacity of diverse attention mechanisms. As shown in Figure 2.
@@
Previous Transformer-based forecasters, such as FEDformer (Zhou et al., 2022) and MultiPatchFormer (Naghashi et al., 2025), tend to overprioritize temporal feature extraction, introducing various complex mechanisms to model temporal dependencies. This design often incurs substantial computational overhead without delivering notable performance gains. We argue that the informational content in the temporal dimension is inherently limited and should not be excessively emphasized. In contrast, HTMformer performs full temporal feature extraction solely within the embedding layer, thereby significantly reducing computational complexity.
@@
iTransformer retains all native Transformer components without modification. Instead, it adopts an inverted input representation, enabling the self-attention model to better capture multivariate correlations and thereby yielding satisfactory performance. This underscores the importance of modeling multivariate correlations in capturing the semantics of time series. In CNN-based predictors, graph convolutional networks (Kipf & Welling, 2017; Yu et al., 2018; Guo et al., 2019) explicitly model correlations among variables by leveraging graph adjacency matrices. This paradigm unifies the temporal dynamics of the data with the multivariate structure of variables within a graph-based model, thereby substantially improving performance in time series tasks. However, such explicit modeling entails significant computational overhead, particularly in Transformer architectures. HTMformer is designed to moderately extract multivariate features and integrate them with temporal information within the embedding layer, yielding semantically enriched embeddings. Furthermore, it employs an inverted input strategy to enhance the efficiency and performance of Transformers and reduce the computational costs.
@@ 3.1 Structure Overview
Motivated by the temporal and multivariate correlations in time series, we propose HTMformer, a general-purpose time series forecasting model. It adopts the encoder-only architecture and incorporates an HTME extractor to effectively capture temporal and multivariate features. As shown in Figure 3, the model includes an HTME extractor, a vanilla Transformer encoder layer, and a projection layer. The HTME extractor is designed to deeply extract temporal and multivariate features. After the time series is processed, the Transformer encoder layer is used to learn the relationships between multiple channels. Finally, a linear model serves as the projection layer. RevIN (Kim et al., 2022) is a normalization technique for time series forecasting, which can help models learn and generalize better. Each data batch is first normalized with RevIN before being fed into the model.
@@
3.2 HTME Extractor
To fully capture informative representations from the raw time series, the HTME extractor incorporates two core components: a temporal feature extractor and a multivariate feature extractor, which enable the independent modeling of two dimensions. The final output is obtained by integrating the outputs of these two components. The HTME extractor takes multivariate time series as input and outputs embedding representations that fuse informative features, as illustrated in Figure 3 (Left).
Figure 3: Overall architecture of HTMformer. The HTME layer comprises two independent modules for temporal and multivariate feature extraction, directly replacing the original embeddings. It also employs inverted input to enable the attention mechanism to capture channels correlations.
3.2.1 Temporal Feature Extractor
In the temporal feature extraction module, the time series is first segmented into patches, with each patch treated as a single input token. Convolutional operations are then utilized to capture short-term temporal patterns within local receptive fields. The convolutional outputs are flattened into a high-dimensional tensor that integrates richer temporal information. Then, a linear projection layer is applied to further integrate local temporal features. The extracted features are subsequently projected along the temporal dimension to model the long-term correlations, and the tensor is reshaped to generate the final representation. Compared with conventional embedding layers, combining convolution with patch-based tokenization enables more accurate modeling of temporal contextual dependencies. By treating each channel as an independent input, interference from other features is reduced, thereby enhancing the model’s capability to capture temporal features effectively.
$N_{1}^{4\times k},\,N_{2}^{4\times k},\,\ldots,\,N_{(L-k)N}^{4\times k}$ $=\operatorname{Conv}\bigl(\operatorname{Patch}(X_{\text{in}})\bigr),$ (2)
$E_{1}^{\text{dim}},\,E_{2}^{\text{dim}},\,\ldots,\,E_{N}^{\text{dim}}$ $=\operatorname{Linear}\bigl(\operatorname{Flatten}\bigl(N_{1}^{4\times k},\,N_{2}^{4\times k},\,\ldots,\,N_{(L-k)N}^{4\times k}\bigr)\bigr),$ (3)
$D_{\text{out}}$ $=\operatorname{Linear}\bigl(E_{1}^{\text{dim}},\,E_{2}^{\text{dim}},\,\ldots,\,E_{N}^{\text{dim}}\bigr).$ (4)
3.2.2 Multivariate Feature Extractor
In the multivariate feature extraction module, the time series is first partitioned into patches to accommodate potential temporal lags. The features of each patch are then flattened to aggregate local information, and a linear layer is applied to model the correlations among input variates. These learned relationships are aggregated across time through temporal convolution. The resulting output is fed into a GRU (Cho et al., 2014) network, whose gating mechanism selectively emphasizes or suppresses different inputs, thereby adaptively modulating the contribution of historical information. Finally, the features fused with multivariate information are expanded along the temporal dimension via a convolution layer to match the output of the temporal module.
$\left(P_{1}^{C_{k}},P_{2}^{C_{k}},\ldots,P_{L-k}^{C_{k}}\right)$ $=\operatorname{Flatten}\bigl(\operatorname{Patch}(X_{\text{in}})\bigr),$ (5)
$M^{N}$ $=\operatorname{GRU}\bigl(\operatorname{Linear}\bigl(P_{1}^{C_{k}},P_{2}^{C_{k}},\ldots,P_{L-k}^{C_{k}}\bigr)\bigr),$ (6)
$V_{\text{out}}$ $=\operatorname{Conv}\bigl(M^{N}\bigr),$ (7)
$Y_{\text{in}}$ $=\alpha D_{\text{out}}+(1-\alpha)V_{\text{out}},$ (8)
@@ 3.3 Encoder and Projection
The Transformer encoder layer is composed of a self-attention module and a position-wise feed-forward network. As temporal features have been fully extracted in the embedding stage, we adopt an inverted input design, which enables the self-attention mechanism to directly model global channel correlations. This inverted structure avoids forced alignment of variables within timestamps and facilitates the concatenation of temporal and multivariate dimensions. Consequently, it enables the attention mechanism to learn more appropriate sequence representations, leading to more accurate predictions. This reduces computational complexity from $\mathcal{O}(L^{2})$ in traditional methods to $\mathcal{O}(N^{2})$.
@@
The self-attention block additionally consists of layer normalization and a position-wise feed-forward network with residual connections. It enables the model to learn cross-channel relationships and yield refined representations for each channel. iTransformer(Liu et al., 2024b) observes that, for numerical time-series forecasting tasks, a linear layer is sufficient to generate predictions, since forecasting can be realized by mapping variate-level representations without resorting to complex decoder architectures. Finally, a linear projection head serves as the decoder, mapping the learned representations to the target dimensionality and yielding the final predictions.
Unlike other models that deeply mine multidimensional correlations (Ying et al., 2021; Lai et al., 2018), our model does not explicitly model the associations between variables. Within HTME, a novel strategy is adopted to obtain high-dimensional representations that integrate multivariate information, which are subsequently mapped onto the temporal dimension. We utilize multivariate features as a complement to temporal features to learn richly informative embeddings, thereby achieving the goal of deeply extracting the hybrid temporal and multivariate characteristics of time series.
@@ 3.4 Complexity Analysis
HTMformer is composed of two primary components: the HTME extraction module and a Transformer encoder. The model’s computational complexity is primarily dictated by the Transformer encoder, as the introduction of the inverted input design results in both time and memory complexity scaling as $\mathcal{O}(N^{2})$. Compared with the state-of-the-art Transformer forecaster PatchTST (Nie et al., 2023) and the vanilla Transformer, HTMformer attains lower complexity because the number of variate tokens is fixed and independent of the lookback length, which confers clear advantages for long-term time series forecasting. Moreover, as the key component, the HTME extraction module is flexibly compatible with Transformer variants integrating efficient attention mechanisms, such as Informer or Reformer, It allows for further reductions in computational and memory costs and broadens the model’s applicability to a wider range of forecasting scenarios (see Appendix C).
@@ 4 Experiments
To evaluate both the effectiveness of the proposed HTME strategy in enhancing Transformer-based forecasters and its capacity for scalability, we apply HTME to various Transformer variants. Notably, HTME consistently and significantly improves the performance of Transformers. Moreover, we conduct a comprehensive evaluation of the proposed HTMformer on diverse time series forecasting and analysis tasks to further validate the advantages of the HTME strategy. Across various key tasks, HTMformer consistently outperforms current state-of-the-art methods. The detailed experimental configurations and implementation details can be found in Appendix D.

### v3

@@ Abstract
Existing predictors primarily focus on modeling the temporal dimension, often achieving superior performance even with channel-independent designs. This is because, due to the inherent limitations of time series data, existing multivariate feature extraction strategies tend to introduce substantial noise and brings additional computational overhead. However, these predictors neglect the inherent multivariate correlations within time series, inevitably leading to an inability to accurately identify complex patterns. To address this issue, we propose a novel multivariate extraction strategy, Hybrid Temporal and Multivariate Embedding (HTME). HTME separates the multivariate feature space for denoising and pattern extraction, and maps the features to the temporal feature space, yielding multidimensional embeddings that convey richer and more meaningful sequence representations. This strategy enables predictors focusing on the temporal dimension fully consider multivariate features with minimal additional computational cost, consistently improving their performance. We construct HTME-versions for eight predictors across three different predictive architectures. In all experiments, the resulting HTMPredictors consistently achieve state-of-the-art performance.
@@ 1 Introduction
Long-term time series forecasting holds significant importance in various fields such as finance and economics (Andersen2006), climate science (Mudelsee2019), healthcare (Zeger2006), geophysics (Gubbins2004), industrial monitoring (Truong2022), etc. The key to time series forecasting lies in capturing temporal dependencies and multivariate correlations, while also meeting the demands of real-time prediction (Lan2022IntraInter; Yu2024Revitalizing).
Due to the inherent constraints of time series, the extraction of multivariate features inevitably introduces significant noise(liu2024itransformer; han2023capacity; li2023revisiting). So existing prediction models, such as PatchTST and DLinear, focusing solely on the temporal dimension have superior performance. We point that this strategy of neglecting multivariate correlations inevitably prevents the model from fully capturing complex time-series patterns. The absence of key patterns inevitably prevents the model from effectively fitting diverse and complex time series, which severely limits its accuracy and robustness. Current hybrid stragetys, such as Crossformer, employing a flawed approach to extracting multivariate features. These models introduce extra, complex modules into their backbone networks, which significantly increases computational overhead. Furthermore, severe noise interference due to a shared feature space, leading to a degradation in accuracy (see Appendix A). Therefore, it is necessary to redesign a strategy for extracting multivariate correlations, enabling these models to reconsider them.
@@
To mitigate this issue, this paper proposes a strategy that hybrid temporal and multivariate features in the embedding (HTME) layer , thereby yielding semantically rich embedding representations to overcome the inherent limitations (see Appendix B).
HTME employs parallel and specialized feature extraction modules to separately capture temporal and multivariate features from the raw input, thereby minimizing interference between them. More importantly, the two feature spaces in HTME are disjoint, which confines the noise exclusively to the multivariate feature space. Then, HTME projects the multivariate features into the temporal feature space. HTME is designed to leverage multivariate features to appropriately supplement the temporal feature space, yielding semantically richer embedding representations. This enables subsequent modules to sufficiently extract both temporal patterns and mutilvariate correlations by focusing on just a single dimension without additional moudels. Rather than mixing or concatenating the feature spaces, HTME suppresses noise within the multivariate space while preserves the integrity and consistency of the temporal feature space. HTME consistently enhances the performance of various predictors focusing on temporal dimension with only negligible computational overhead.
Leveraging this design, we construct a novel Transformer-based framework(HTMformers), as illustrated in Figure 2. HTMformers consistently enhance the modeling capabilities of various attention mechanisms. By applying the HTME strategy, even simple models can achieve performance comparable to complex models. HTMformer using Full-Attention as a presentative of HTMformers achieves state-of-the-art performance on real-world benchmarks, as shown in Figure 1.
We further applied the HTME strategy to other forecasting models with different architectures, and it consistently improved their performance.This demonstrates the validity and strong generalizability of the HTME strategy.
The contributions of this work are as follows:
  • We propose the Hybrid Temporal and Multivariate Embedding (HTME) strategy. It enables predictors focusing on the temporal dimension to fully account for multivariate features, while introducing only minimal computational overhead.
  • We introduce a novel forecasting framework, HTMformers, which consistently improves the performance of attention mechanisms in forecasting. The representation HTMformer consistently achieves state-of-the-art performance across various benchmarks.
  • We selected seven widely-used forecasting models and developed their HTME-augmented versions. Extensive experiments demonstrate that HTME consistently enhances the ability of diverse architectures to model complex time series, particularly for Transformer framework.
@@
Previous time series forecasters, especially Transformer-based, such as FEDformer (zhou2022fedformer) and MultiPatchFormer (naghashi2025multipatchformer), tend to overprioritize temporal feature extraction. This design often incurs significant computational overhead and can lead to model overfitting. In contrast, HTME performs a deep embedding of temporal features at the embedding layer and augments them with multivariate features. This process generates semantically richer embedding representations, enabling various models to capture finer-grained features and achieve superior performance.
@@
GCN (kipf2017gcn; yu2018stgcn; guo2019astgcn) extracts correlations among variables by leveraging graph adjacency matrices, while iTransformer (liu2024itransformer) extracts multivariate correlations using Attention. Their success highlights the importance of multivariate correlations. However, this approach of explicitly modeling multivariate correlations incurs significant computational overhead. In contrast, the HTME strategy employs a lightweight module to independently extract multivariate features without introducing additional complex modules.
Channel Dependence strategy.
Recent studies (han2023capacity; li2023revisiting) have found that Channel Dependence model makes it prone to capturing spurious correlations from noise and overemphasizing coincidental patterns in the training data. This is because existing models, such as Crossformer, make the multivariate features and temporal features share a single space. HTME independently extracts multivariate features and projects them to the temporal feature space, allowing subsequent modules to remain focused on the temporal dimension, which significantly suppresses noise.
@@ 3 HTMformer
3.1 HTME Extractor
To fully capture informative representations from the raw time series, the HTME extractor incorporates two novel components: a temporal feature extractor that can more effectively extract temporal patterns and a lightweight multivariate feature extractor.This parallel design which enable the independent modeling reduces interference between features. Finally, we employ a weighted summation for feature fusion, enabling multivariate features to complement temporal features, thereby yielding a richer representation. The lightweight design ensures minimal additional computational overhead.
Within the two branches, we adopting a bottom-up fusion followed by a top-down decomposition strategy. This strategy encourages the model to focus on specific dimension, thereby avoiding interference between dimensions, while guiding the model to emphasize import patterns, resulting in more effective feature embeddings.
The HTME extractor takes raw multivariate time series $X_{\text{in}}\in\mathbb{R}^{L\times N}$ as input and outputs embedding representations that fuse informative features, as illustrated in Figure 3.
3.1.1 Temporal Feature Extractor
The key to temporal modeling is to extract both short-term and long-term dependencies. It is worth noting that short-term correlations are predominant, while the influence of long-term correlations should not be neglected. The strategy of first merging and then progressively decomposing features effectively guides the model to focus step-by-step on the most important patterns.
First, we adopt a patching strategy to segment short-term temporal representations from the original sequence, and iteratively merge all channels to obtain a short-step representation $T^{N\times(L-K),K}$ .It is worth noting that the process merely partitions the time series without introducing trainable parameters at this stage.
$P_{1},P_{2},P_{3},\dots,P_{L-K}$ $=X_{1:1+K},X_{2+K},X_{3+K},\dots,X_{L-K:L}$ (1)
$T^{N\times(L-K),K}$ $=\text{Stack}(P_{1},P_{2},P_{3},\dots,P_{L-K})$ (2)
Short-term dependencies are particularly important. this module first prioritizes the deep extraction of short-term temporal features. Specifically, we apply multiple convolutional filters to capture short-term temporal patterns at multiple scales. We then fuse the convolutional outputs via a linear layer to obtain a feature representation enriched with short-term dependencies.
$E^{N\times(L-K),8}$ $=\text{Linear}(\text{Conv}(T^{N\times(L-K),1,K}))$ (3)
Finally, we unfold the long temporal dimension and employ a second linear layer to capture the long-term dependencies within the sequence, yielding a deep embedding that encapsulates both short-term and long-term temporal dependencies.
$D_{\text{out}}$ $=\text{Linear}(\text{Flatten}(E^{N\times(L-K),8}))$ (4)
Compared with existing embedding strategies, our method achieves a more fine-grained modeling of temporal contextual dependencies using only one convolutional layer and two linear layers.
Figure 3: HTME consists of two independent and parallel feature extraction modules. It employs a strategy of first merging the outputs and then hierarchically decomposing the resulting representation.
3.1.2 Multivariate Feature Extractor
We adopt the same merge‑decompose strategy as described above. We segment and stack the time series to obtain $T^{N\times(L-K),K}$, following exactly the same procedure as in the aforementioned module.
Subsequently, we merge the short‑term temporal dimension with the multivariate dimension, and employ a linear layer to process the multi‑step variables, thereby obtaining multivariate features $N^{(L-K),128}$ that account for short‑term temporal lags.
$N^{(L-K),128}$ $=\text{Linear}(T^{(L-K),K\times N})$ (5)
Similar to Eq. (4), we employ multiple convolutional filters to extract multivariate features at different scales, resulting in a more fine-grained multi-scale multivariate feature block. As this module is substantial in size, we therefore apply a specialized GRU network to aggregate the multi-scale features and obtain the integrated representation $B^{N}$.
$B^{N}$ $=\text{GRU}(\text{Conv}(N^{(L-K),128}))$ (6)
Considering that the correlations among multivariate variables may differ across time stamps, we apply different convolutional kernels to project the integrated representation $B^{N}$ onto the temporal dimension, thereby obtaining the final representation.
$V_{\text{out}}$ $=\text{Conv}(B^{N})$ (7)
$Y_{\text{in}}$ $=\alpha D_{\text{out}}+(1-\alpha)V_{\text{out}},$ (8)
@@ 3.2 Structure Overview
We construct HTMformer based on iTransformer, a state-of-the-art and generic Transformer architecture. It keeps the native components of the original Transformer intact. This implies that other components within HTMformer can be flexibly interchanged with their respective variants. As shown in Figure 4, the model includes an HTME extractor, a vanilla Transformer encoder layer, and a projection layer. RevIN (kim2022reversible) is a normalization technique, which can help models learn and generalize better. Each data batch is first normalized with RevIN before being fed into the model.
@@
Figure 4: Architecture of HTMformer. The HTME directly replacing the iTransformer embeddings. It also employs inverted input to enable the attention mechanism to capture channels correlations.
@@ 3.3 Encoder and Projection
The Transformer encoder layer is composed of a self-attention module and a position-wise feed-forward network. We adopt an inverted input design, which enables the self-attention mechanism to directly model global channel correlations. This inverted structure avoids forced alignment of variables within timestamps and facilitates the concatenation of temporal and multivariate dimensions. Consequently, it enables the attention mechanism to learn more appropriate sequence representations.
@@
The self-attention block additionally consists of layer normalization and a position-wise feed-forward network. Finally, a linear projection head serves as the decoder, mapping the learned representations to the target dimensionality and yielding the final predictions. iTransformer (liu2024itransformer) observes that this Transformer‑based predictor architecture is more efficient for time series forecasting tasks.
@@ 3.4 Complexity Analysis
HTME consists of two parallel extraction modules. The time and space complexities of the Feature Extraction Modules are both O(LN). For each dimension, the computational complexity is linear, which guarantees the high efficiency and scalability of HTME. The HTME module does not introduce significant additional computational overhead. The primary computational cost of HTMProdectors stems from the underlying base predictor they are built upon.
@@ 4 Experiments
To validate the effectiveness and transferability of the HTME strategy, we integrated it into diverse forecasting backbones and evaluated its impact on predictive performance. Moreover, we conducted a comprehensive evaluation of the proposed HTMformer on diverse forecasting tasks to further validated the advantages of the HTME strategy. We also evaluated the computational overhead of HTME, showing that it introduces only modest additional cost. The detailed experimental configurations and implementation details can be found in Appendix D.